### v1

@@ 2.1 The original Forward-Forward algorithm
Predefined positive and negative examples are successively presented to the network’s input. The possibility of a positive example $\boldsymbol{x}_{i}$ being recognized as positive and a negative example $\boldsymbol{x}_{j}$ being recognized as negative by the network are defined as $p_{\text{pos}}(\boldsymbol{x}_{i})=\sigma(G_{i}^{(l)}-\Theta_{\text{pos}}^{(l)})$ and $p_{\text{neg}}(\boldsymbol{x}_{j})=\sigma(\Theta_{\text{neg}}^{(l)}-G_{j}^{(l)})$ respectively. The sigmoid function $\sigma(x)=\frac{1}{1+e^{-x}}$ evaluates the effectiveness of the separation, where $\Theta_{\text{pos}}^{(l)}$ and $\Theta_{\text{neg}}^{(l)}$ are fixed values that serve as hyperparameters of the network.
@@
Inspired by Noise-Contrastive Estimation [48], which aims to distinguish positive examples from noisy examples, the objective of FF learning is to increase the goodness score for positive input examples so that it significantly exceeds the threshold ($p_{\text{pos}}(\boldsymbol{x}_{i})\rightarrow$ 1) and to decrease the goodness score for negative input examples so that it falls well below the threshold ($p_{\text{neg}}(\boldsymbol{x}_{j})\rightarrow$ 1). Weight updates are performed locally by combining the gradients derived from both positive and negative examples:
@@
Hinton [19] proposed that FF can be used for self-supervised representation learning by using real data vectors as the positive examples and engineered data vectors as the negative examples. The negative examples should be carefully designed, rather than relying on random corruption methods of the training data like noise injection or occlusion. The negative data needs to be sufficiently challenging to ensure the network learns useful information. To make FF focus on the long-range correlations in images that characterize shapes, the negative data should have very different long-range correlations while maintaining similar short-range correlations. For the MNIST dataset, this can be achieved by creating a mask with large regions of ones and zeros. As proposed in the original framework [19], negative data is then generated by combining one digit image multiplied by the mask with a different digit image multiplied by the inverse of the mask, as illustrated in Fig 2a. Although this method performs well for MNIST (as shown in the Unsupervised panel in Fig 1), it does not easily translate to more complex image databases like CIFAR-10 and STL-10, resulting in limited accuracy on these benchmarks.
2.2 Constrastive self-supervised learning
In this article, we draw inspiration from contrastive self-supervised learning methods to construct positive and negative examples for FF, applicable to any image database and beyond. Contrastive learning, illustrated in Fig 2c, is a self-supervised technique designed to learn representations by contrasting positive pairs against negative pairs. This approach relies heavily on data augmentation, where a positive pair consists of two different augmented views of the same data sample, and negative pairs are constructed from different samples. While FF shares strong similarities with contrastive learning, it also has key differences.
Both methods utilize the concept of contrasting positive and negative examples to guide the learning process. They also both aim to learn meaningful data representations without labeled data in the first phase, allowing to extract information from the hidden layers in the second phase by training a linear classifier.
However, their loss functions are fundamentally different. FF focuses on the absolute goodness of positive versus negative examples, decoupling them in the loss. On the other hand, general contrastive losses employ relative comparisons through distances or similarities between positive and negative examples.
Additionally, the neural network optimization method differs. FF defines a local loss function at each layer and avoids backpropagation through hidden layers, whereas typical contrastive learning approaches train deep networks end-to-end. Lastly, self-supervised FF emphasizes data corruption to generate negative examples, focusing on specific characteristics of corruption to enhance learning, while contrastive learning relies heavily on varied data augmentation techniques.
3 Self-Constrastive Forward-Forward algorithm and results
We draw inspiration from contrastive self-supervised learning algorithms to propose the Self-Contrastive Forward-Forward (SCFF) algorithm, which adresses the drawbacks of FF, including the complexity of generating negative examples and its inability to generalize well to convolutional neural networks in a purely forward manner.
3.1 Creating the negative and positive examples
Instead of contrasting the representations in the feature space as in contrastive self-supervised learning (Fig 2c), SCFF directly takes pairs of positive and negative images as inputs to the neural network (Fig 2d). More specifically, given a batch of $N$ training examples, and for a randomly selected example $\boldsymbol{x}_{k}$ ($k\in\{1,N\}$) in the batch, the positive example $\boldsymbol{x}_{i,\text{pos}}^{(0)}$ (the number 0 is the layer index) is the concatenation of two repeated $\boldsymbol{x}_{k}$, i.e., $\boldsymbol{x}_{i,\text{pos}}^{(0)}=[\boldsymbol{x}_{k},\boldsymbol{x}_{k}]^{T}$. The negative example $\boldsymbol{x}_{j,\text{neg}}^{(0)}$ is obtained by concatenating $\boldsymbol{x}_{k}$ with another example $\boldsymbol{x}_{n}$ ($n\neq k$) in the batch, i.e., $\boldsymbol{x}_{j,\text{neg}}^{(0)}=[\boldsymbol{x}_{k},\boldsymbol{x}_{n}]^{T}$ (or $[\boldsymbol{x}_{n},\boldsymbol{x}_{k}]^{T}$ ). Fig 3a shows some instances of generated positive and negative examples from the original training batch for the STL-10 dataset.
Considering the case of a fully connected network, the concatenated pair of images (positive or negative examples) are sent as inputs to the network. The outputs for the positive and negative examples from the first layer can be written respectively as $\boldsymbol{y}_{i,\text{pos}}^{(0)}=f(W_{1}\boldsymbol{x}_{k}+W_{2}\boldsymbol{x}_{k})$, $\boldsymbol{y}_{j,\text{neg}}^{(0)}=f(W_{1}\boldsymbol{x}_{k}+W_{2}\boldsymbol{x}_{n})$, where $f$ is the ReLU activation function. The weight matrices $W_{1}$ and $W_{2}$ correspond to the connections to the two images. In practice, we set $W_{1}=W_{2}$ because the gradient of the loss function with respect to $W_{1}$ and $W_{2}$ converges to the same value. Setting $W_{1}=W_{2}$ accelerates the training speed and improves the performance. Intuitively, this can be understood by recognizing that swapping the positions of $\boldsymbol{x}_{k}$ and $\boldsymbol{x}_{n}$ in the concatenated image should not affect the output neural activities. For a rigorous mathematical proof, see Appendix A.
Figure 4: Probability distributions of relative positions between positive and negative examples. a Theoretical distributions of positive examples from two different classes with distinct means ($02\mu_{1}=$ 0 and $152\mu_{2}=$ 15) and identical variance ($42\Sigma=$ 4) are shown with blue and orange curves, respectively. The theoretical distribution of negative examples derived from the two classes using the formula 3.1 is depicted by the grey curve. b Continuous probability density of LDA applied to the IRIS dataset, displaying contours for positive examples in green, red, and blue, and for negative examples in grey.
The concept of SCFF can be understood through the lens of Noise Contrastive Estimation (NCE). In NCE, a key insight is that “the noise distribution should be close to the data distribution, because otherwise, the classification problem might be too easy and would not require the system to learn much about the structure of the data”[48]. Our method of generating the positive and negative examples aligns with this principle if we treat the negative examples as “noise data”. We assume that the data samples for each class follow a multivariate Gaussian distribution with a shared covariance matrix $\Sigma$ and that each class is statistically independent of the others—assumptions commonly employed in various statistical models [50]. Furthermore, noting that the input weight matrices $W_{1}$ and $W_{2}$ are identical, i.e. $W=W_{1}=W_{2}$, the concatenated inputs processed by the network are simplified as follows: the positive example becomes $\boldsymbol{y}_{i,\text{pos}}^{(0)}=f(W_{1}\boldsymbol{x}_{k}+W_{2}\boldsymbol{x}_{k})=f(W(2\boldsymbol{x}_{k}))$, and the negative example is $\boldsymbol{y}_{j,\text{neg}}^{(0)}=f(W_{1}\boldsymbol{x}_{k}+W_{2}\boldsymbol{x}_{n})=f(W(\boldsymbol{x}_{k}+\boldsymbol{x}_{n}))$. This is equivalent to treating the positive examples as $\boldsymbol{x}_{i,\text{pos}}^{(0)}=2\boldsymbol{x}_{k}$ and negative examples as $\boldsymbol{x}_{j,\text{neg}}^{(0)}=\boldsymbol{x}_{k}+\boldsymbol{x}_{n}$. Therefore, the distributions of positive examples $\boldsymbol{x}_{i,\text{pos}}^{(0)}$ and negative examples $\boldsymbol{x}_{j,\text{neg}}^{(0)}$ follow :
where $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$ are means of two different classes respectively. Theoretically, the negative examples always lie somewhere between two different clusters of positive examples in the sample space, as illustrated in Fig. 4a for the one-dimensional case. For practical analysis with a real-world dataset, we visualized the distributions of positive and negative examples from the IRIS dataset [51] using 2D linear discriminant analysis (LDA), which distinguishes between three different types of irises, as shown in Fig.4b. This visualization shows that the negative examples are positioned between different clusters of positive examples, suggesting that they contribute to pushing apart and further separating adjacent positive examples as they are mapped into higher-dimensional space during training. Additionally, negative examples are formed by combining two examples from different classes, enriching the diversity of negative examples and leading to more robust training. For a detailed analysis of how the LDA components evolve during training as the input data is mapped into the feature space and more theoretical results, please refer to Appendix B.
3.2 Training procedure
@@
We evaluate SCFF on different image datasets including MNIST [45], CIFAR-10 [46] and STL-10 [47] (results in Fig. 1 and Fig. 5), as well as an audio dataset Free Spoken Digit Dataset (FSDD) [52] (results in Fig. 6).
Each layer of the network was fully trained and frozen before training the next one. After unsupervised training with SCFF, we froze the network and trained a linear downstream classifier [53, 9] with the back-propagation method on representations created by the network using the labeled data. The linear classifier was optimized using cross-entropy loss. The accuracy of this classification serves as a measure of the quality of the learned representations. To ensure successful training and high accuracy in the final model, several important techniques were utilized during the training process. These techniques include the “Triangle” method [47, 11] for transmitting information between layers, adding a penalty term in the loss function to ensure stable training, and applying an extra pooling layer to retrieve information at each layer. For further details, see the Methods section.
All details about the impact of hyperparameters and the training of the linear classifier are provided in Appendix F and G.
3.3 Fully connected network: MNIST
@@
On the MNIST dataset, when trained on a two-layer fully connected network with 2000 hidden neurons each, SCFF achieves a test precision of 98.7%, which is comparable to the performance achieved by backpropagation [19]. This surpasses previously published benchmarks on other biologically inspired algorithms, such as 97.9% in [9], 98.42% in [8] (supervised training), and 96.6% in [7]. The full comparison is shown in the right column of the green frame in Fig 1.
3.4 Convolutional neural networks: CIFAR-10 and STL-10
1 layer2 layers3 layers505050506060606070707070808080809090909072.172.172.172.178.378.378.378.380.880.880.880.874.474.474.474.483.883.883.883.884.984.984.984.9Accuracy (%)a CIFAR-10SCFFBackprop
@@ 3.4 Convolutional neural networks: CIFAR-10 and STL-10
The convolutional neural network processes three-dimensional color images. The original images are concatenated along the channel dimension to form positive or negative inputs (see Fig 3). The output of each convolutional layer is represented as a three-dimensional vector $\boldsymbol{y}_{i,\text{pos}}^{(l)}$ (or $\boldsymbol{y}_{i,\text{neg}}^{(l)}$) $\in\mathbb{R}^{C\times H\times W}$. The Loss function at layer $l$ is then defined as:
@@
For the CIFAR-10 and STL-10 datasets, we employed convolutional neural networks with architectures identical to those in [9] to extract features. SCFF achieves a test accuracy of 80.75% with a three-layer convolutional architecture (number of filters each layer: 96-384-1536) on CIFAR-10 and 77.3% with a four-layer convolutional architecture (number of filters each layer: 96-384-1536-6144) on STL-10. These results surpass the previous state-of-the-art accuracies for purely-forward unsupervised learning, of 80.3% on CIFAR-10 and 76.2% on STL-10 achieved using the SoftHebb algorithm [9]. This demonstrates the significant potential of the SCFF method to scale effectively to more complex datasets and architectures. The full comparison is shown in the left column of the green frame in Fig. 1.
@@
We also compared the test accuracies at each layer using SCFF and Backpropagation (BP) methods on CIFAR-10 and STL-10, as shown in 5. Notably, for STL-10, SCFF achieved a final layer performance of 77.3% higher than the one of BP: 77.02%(Fig. 5b). This is because the STL-10 dataset contains a large amount of unlabeled images, which limits the effectiveness of supervised BP training. By fine-tuning SCFF with end-to-end BP on the few labelled STL-10 examples, SCFF’s accuracy further improves to 80.13%. This demonstrates that SCFF is highly suitable for unsupervised pretraining followed by supervised BP training, making it ideal for semi-supervised learning approaches.
@@
Unlike other unsupervised learning methods, where the result is obtained solely from the final layer’s output, SCFF integrates neuron information with the linear classifier from intermediate layers, leading to more comprehensive feature extraction [19]. For CIFAR-10 (Fig. 5 a), the test accuracy for the two-layer and three-layer models was obtained by combining the outputs of all previous layers (pooled information for dimensionality reduction; see Methods section) before feeding them into the final linear classifier. However, because the STL-10 dataset consists of high-quality images, the number of output neurons in each layer is much larger than that in CIFAR-10. Therefore, for the STL-10 dataset, we did not combine outputs from previous layers for training the linear classifier, with the exception of the fourth layer. In this case, we combined the outputs from both the third and fourth layers for the final classification, resulting in a 1% improvement in accuracy compared to using only the fourth layer’s outputs as input to the linear classifier.
@@
3.5 Recurrent neural network: FSDD audio dataset
The original FF paper [19] introduces a specialized recurrent neural network (RNN) that models top-down effects using repeated static images as input for each time frame. In contrast, our work focuses on training an RNN that processes time-varying inputs.
@@ 3.5 Recurrent neural network: FSDD audio dataset
We employ the Free Spoken Digit Dataset (FSDD), a standard benchmark task for evaluating RNN training performance [55, 56, 57]. The FSDD is a collection of audio recordings where speakers pronounce digits (0-9) in English. We follow the standard procedure consisting in extracting frequency domain information at different time intervals, here through Mel-Frequency Cepstral Coefficient (MFCC) features (39 channels) [58]. Plots of the evolution of MFCC feature with time are shown in Fig. 6 for the digits 3 and 8. The SCFF method forms positive and negative examples by concatenating the same input for positive examples, and different ones for negative examples. Fig. 6a presents a negative example which is generated by concatenating MFCC features from two different digits. The goal of the task is to recognize the digit after feeding in the full sequence, from the output of the network at the last time step.
@@
We train a Bi-directional Recurrent Neural Network (Bi-RNN) in an unsupervised way using the SCFF method to classify the digits. The procedure that we use for this purpose is illustrated in Fig. 6a. Unlike conventional uni-directional RNNs, where sequential input is processed step by step in a single direction, resulting in a sequence of hidden states from $H_{0}$ to $H_{T}$ (as depicted in the bottom RNN in Fig. 6a), the Bi-RNN comprises two RNNs that process the input in parallel in both forward and backward directions. This results in hidden states evolving from $H_{0}$ to $H_{T}$ in the forward RNN and from $H_{T}^{*}$ to $H_{0}^{*}$ in the backward RNN, as shown in the top portion of the figure. The red regions in the figure highlight the features at different time steps. This bidirectional structure is particularly advantageous for tasks where context from both preceding and succeeding audio frames is critical, such as speech recognition, enhancing model performance compared to conventional uni-directional RNNs [59].
@@
Figure 6: Bi-directional RNN results on FSDD dataset. a. Training procedure of SCFF on a Bi-RNN. In the first stage, unsupervised training is performed on the hidden connections (both input-to-hidden and hidden-to-hidden transformations) using positive and negative examples. Positive examples are created by concatenating two identical MFCC feature vectors of a digit along the feature dimension, while negative examples are generated by concatenating MFCCs from two different digits, as illustrated in the figure. At each time step, the features are sequentially fed into the Bi-RNN (RNN and RNN). The red regions indicate features at different time steps. In the second stage, a linear classifier is trained using the final hidden states from both RNNs, i.e., $H_{T}$ and $H_{0}^{*}$ as inputs for classification task. b. Comparison of test accuracy for the linear classifier trained on Bi-RNN outputs. The yellow curve represents accuracy with untrained (random) hidden neuron connections, the blue curve shows results from training with SCFF, the green curve shows Backprop results
@@
After the first stage of unsupervised training, a linear classifier is trained on the hidden states from the final time step in both directions, as shown in the bottom of the Fig. 6a. The blue, orange and green curves in Fig. 6b depict the test accuracy of the linear output classifier with hidden connections trained using SCFF, with random (untrained) hidden connections, and with Backpropagation methods, respectively. SCFF achieves a test accuracy of 90.3% when trained with a one-layer Bi-RNN containing 500 hidden neurons in each direction (refer to Appendix E for further architectural details). It is below the performance of BackPropagation Though Time, reaching 96% accuracy on this small task, but well above the model with untrained (random) hidden connections which plateaus at 80.7%.
This result constitutes the first successful application of the FF approach to sequential data in an unsupervised manner. Unlike the BPTT method for training RNNs, SCFF avoids the issues of vanishing and exploding gradients, as the gradients at each time step are calculated independently. This eliminates the dependency between time steps, providing a more stable training process.
@@
It is also interesting to note that the network with untrained hidden and input connections is akin to Reservoir Computing, a method that is often used to leverage physical systems on sequential data for neuromorphic applications [60]. SCFF provides a way to train these input and hidden layer connections in a simple, hardware-compatible way, and opens the path to a considerable gain of accuracy. This achievement opens the door for its extension to more complex tasks involving temporal sequences and its potential use in neuromorphic computing domains, such as dynamic vision sensors [61].

### v2

@@ 2.2 Training procedure
  • Loss Function: Predefined positive and negative examples are successively presented to the network’s input. The possibility of a positive example $\boldsymbol{x}_{i}$ being recognized as positive and a negative example $\boldsymbol{x}_{j}$ being recognized as negative by the network are defined as $p_{\text{pos}}(\boldsymbol{x}_{i})=\sigma(G_{i}^{(l)}-\Theta_{\text{pos}}^{(l)})$ and $p_{\text{neg}}(\boldsymbol{x}_{j})=\sigma(\Theta_{\text{neg}}^{(l)}-G_{j}^{(l)})$ respectively. The sigmoid function $\sigma(x)=\frac{1}{1+e^{-x}}$ evaluates the effectiveness of the separation, where $\Theta_{\text{pos}}^{(l)}$ and $\Theta_{\text{neg}}^{(l)}$ are fixed values that serve as hyperparameters of the network.
@@
    The loss function encourages the network to increase the goodness for positive examples so that it significantly exceeds the threshold ($p_{\text{pos}}(\boldsymbol{x}_{i})\rightarrow$ 1) and to decrease the goodness score for negative input examples so that it falls well below the threshold ($p_{\text{neg}}(\boldsymbol{x}_{j})\rightarrow$ 1). At layer $l$, the loss function is defined as:
@@
  • Normalization and Standardization
    To ensure stability during training, SCFF applies two key pre-processing steps:
    • Dataset-wide Normalization: Each input image is normalized by subtracting the mean and dividing by the standard deviation, calculated per channel across the entire dataset.
    • Per-image Standardization: Each individual image is standardized so that its pixel values have a mean of 0 and a standard deviation of 1. This ensures consistent input scaling, which is particularly important for unsupervised learning.
    Other techniques include the “Triangle” method [54, 57] for transmitting information between layers, adding a penalty term in the loss function to ensure stable training, and applying an extra pooling layer to retrieve information at each layer. For further details, see the Methods section. All details about the impact of hyperparameters and the training of the linear classifier are provided in Appendix G and H.
3 Results
@@
We evaluate SCFF on different image datasets including MNIST [51], CIFAR-10 [52], Tiny ImageNet [53] and STL-10 [54] (results in Table 1 and Fig. 3), as well as an audio dataset Free Spoken Digit Dataset (FSDD) [58] (results in Fig. 4). Across all benchmarks, SCFF surpasses all state-of-the-art algorithms that combine a local learning rule with fully-forward operation, while maintaining its advantage in hardware adaptability.
3.1 Multilayer Perceptron (MLP): MNIST
@@
On the MNIST dataset, SCFF achieves a test accuracy of 98.70% $\pm$ 0.01% when trained on a two-layer fully-connected network (MLP) with 2000 hidden neurons per layer, which is comparable to the performance achieved by backpropagation [16]. This surpasses previously published benchmarks on other biologically-inspired algorithms applied to MLPs, including 97.8% in [59], 98.42% in [60] (supervised training), and 96.6% in [61]. The full comparisons are presented in Table 2, 3, and 5.
3.2 Convolutional Neural Networks (CNN): MNIST, CIFAR-10 and Tiny ImageNet
@@
The convolutional neural network (CNN) processes three-dimensional color images. The original images are concatenated along the channel dimension to form positive or negative inputs (see Fig. 2). The output of each convolutional layer is represented as a three-dimensional vector $\boldsymbol{y}_{i,\text{pos}}^{(l)}$ (or $\boldsymbol{y}_{i,\text{neg}}^{(l)}$) $\in\mathbb{R}^{C\times H\times W}$. The Loss function at layer $l$ is then defined as:
@@
We evaluate SCFF using a three-layer CNN with the same architecture as the state-of-the-art SoftHebb, using 96, 384, and 1536 filters respectively [11]. On MNIST, our results show that SCFF achieves a test accuracy of 99.37% $\pm$ 0.06% when trained with a CNN, comparable to the 99.35% achieved by SoftHebb [11].
1 layer2 layers3 layers505050506060606070707070808080809090909072.172.172.172.178.378.378.378.380.880.880.880.874.574.574.574.584.184.184.184.184.984.984.984.9Accuracy (%)a CIFAR-10SCFFBackprop
For the CIFAR-10 dataset, SCFF achieves an accuracy of 80.75% $\pm$ 0.12%, surpassing the previous state-of-the-art accuracies for purely-forward unsupervised learning, of 80.3% on CIFAR-10 achieved using the SoftHebb algorithm [11].
We also compared the test accuracies at each layer using SCFF and Backpropagation (BP) methods, as shown in Fig. 3a. The comparison reveals that SCFF can effectively capture complex representations at deeper layers, similar to BP. Additionally, the evolution of accuracy with more layers indicates the scalability of SCFF and its robustness in deeper architectures.
To further evaluate SCFF on more complex datasets, we apply an AlexNet-inspired architecture with five convolutional layers (filter sizes: 64-192-384-256-256) to Tiny ImageNet, which is a subset of ImageNet [62] containing 200 object classes. SCFF achieves a top-1 accuracy of 35.67% $\pm$ 0.42% and a top-5 accuracy of 59.75% $\pm$ 0.18%, significantly outperforming previous attempts to apply biologically inspired local learning algorithms to this database. Specifically, SCFF surpasses the Hebbian learning-based approach of Ref. [12], which achieved a top-5 accuracy of 36.99%, and the Inference Learning Algorithm, which achieved a top-5 accuracy of 47.53% [63]. These results highlight SCFF’s ability to learn hierarchical representations efficiently in larger-scale image classification tasks.
3.3 STL-10: Semi-Supervised Learning
The STL-10 dataset is designed for semi-supervised learning tasks, where the majority of the data is unlabeled. We train SCFF using a four-layer convolutional neural network (CNN) with 96, 384, 1536 and 6144 filters respectively [11]. We compare its performance against both traditional supervised learning methods and other local learning algorithms.
@@ 3.3 STL-10: Semi-Supervised Learning
Notably, for STL-10, SCFF achieved a final layer performance of 77.30% $\pm$ 0.12%, higher than the one of BP: 77.02% $\pm$ 0.22% (Fig. 3b). This is because the STL-10 dataset contains a large amount of unlabeled images, which limits the effectiveness of supervised BP training. By fine-tuning SCFF with end-to-end BP on the few labelled STL-10 examples, SCFF’s accuracy further improves to 80.13%. This demonstrates that SCFF is highly suitable for unsupervised pretraining followed by supervised BP training, making it ideal for semi-supervised learning approaches.
@@
Unlike other unsupervised learning methods, where the result is obtained solely from the final layer’s output, SCFF integrates neuron information with the linear classifier from intermediate layers, leading to more comprehensive feature extraction [16]. For CIFAR-10 (Fig. 3a), the test accuracy for the two-layer and three-layer models was obtained by combining the outputs of all previous layers (pooled information for dimensionality reduction; see Methods section) before feeding them into the final linear classifier. For the STL-10 dataset, we combined the outputs from both the third and fourth layers for the final classification, resulting in a 1% improvement in accuracy compared to using only the fourth layer’s outputs as input to the linear classifier.
@@
3.4 Comparison of Greedy Layer-wise Training and Joint Training
We compared the results of greedy layer-wise training and joint training on multiple datasets. Our results indicate that both approaches yield comparable performance. For instance, on CIFAR-10, joint training reaches an accuracy of 80.60% $\pm$ 0.15%, whereas layer-wise training achieves 80.75% $\pm$ 0.12%. A similar trend is observed on STL-10, where joint training attains 77.14% $\pm$ 0.04%, while layer-wise training achieves 77.30% $\pm$ 0.14%. These findings indicate that layer-wise training does not degrade feature representations, as confirmed by stable accuracy trends across deeper layers.
The similarity in performance between these two training strategies is due to the fact that weight updates in earlier layers do not affect the normalized inputs passed to subsequent layers [16]. Specifically, all activations in a given layer scale by the same factor after an update, and layer normalization cancels out this scaling effect. This ensures that greedy layer-wise training prevents error accumulation, maintains stability, and yields feature representations comparable to joint training. Full derivations can be found in Appendix C.
Beyond accuracy, layer-wise training offers practical benefits: it enables more efficient hardware implementation and reduces memory constraints—making it particularly advantageous for neuromorphic computing. Additionally, training each layer independently allows for incremental learning and adaptability, making SCFF more flexible in scenarios where training resources are limited or where continual learning is desired.
3.5 Free Spoken Digit Dataset (FSDD): Sequential Data
Prior research in forward-only learning, including Fast Weights [65], differentiable plasticity [66], short-term plasticity [67, 68], and e-prop [28], has demonstrated the viability of forward-only approaches for sequence learning. While the original FF paper [16] includes a multi-layer recurrent neural network, it uses a static MNIST image repeated over time frames as input, with the objective of modeling top-down effects. Another implementation demonstrates a limited form of sequence learning with a fully connected network, but this architecture could not handle real-time sequential data due to the absence of recurrence. As a result, FF has yet to be extended to effectively handle recurrent network scenarios for time-varying inputs. One of SCFF’s most significant advancements over FF is its ability to handle time-series data.
@@ 3.5 Free Spoken Digit Dataset (FSDD): Sequential Data
We employ the Free Spoken Digit Dataset (FSDD), a standard benchmark task for evaluating RNN training performance [69, 70, 71]. The FSDD is a collection of audio recordings where speakers pronounce digits from 0 to 9 in English. We follow the standard procedure that consist in extracting frequency domain information at different time intervals, here through Mel-Frequency Cepstral Coefficient (MFCC) features with 39 channels [72]. Plots of the evolution of MFCC features with time are shown in Fig. 4 for the digits 3 and 8. The SCFF method forms positive and negative examples by concatenating the same input for positive examples, and different ones for negative examples. Fig. 4a presents a negative example which is generated by concatenating MFCC features from two different digits. The goal of the task is to recognize the digit after feeding in the full sequence, from the output of the network at the last time step.
@@
We train a Bi-directional Recurrent Neural Network (Bi-RNN) in an unsupervised way using the SCFF method to classify the digits. The procedure that we use for this purpose is illustrated in Fig. 4a. Unlike conventional uni-directional RNNs, where the sequential input is processed step by step in a single direction, resulting in a sequence of hidden states from $H_{0}$ to $H_{T}$ (as depicted in the bottom RNN in Fig. 4a), the Bi-RNN comprises two RNNs that process the input in parallel in both forward and backward directions. This results in hidden states evolving from $H_{0}$ to $H_{T}$ in the forward RNN and from $H_{T}^{*}$ to $H_{0}^{*}$ in the backward RNN, as shown in the top portion of the figure. The red regions in the figure highlight the features at different time steps. This bidirectional structure is particularly advantageous for tasks where context from both preceding and succeeding audio frames is critical, such as speech recognition, enhancing model performance compared to conventional uni-directional RNNs [73].
@@
Figure 4: Bi-directional RNN results on FSDD dataset. a. Training procedure of SCFF on a Bi-RNN. In the first stage, unsupervised training is performed on the hidden connections (both input-to-hidden and hidden-to-hidden transformations) using positive and negative examples. Positive examples are created by concatenating two identical MFCC feature vectors of a digit along the feature dimension, while negative examples are generated by concatenating MFCCs from two different digits, as illustrated in the figure. At each time step, the features are sequentially fed into the Bi-RNN (RNN and RNN). The red regions indicate features at different time steps. In the second stage, a linear classifier is trained using the final hidden states from both RNNs, i.e., $H_{T}$ and $H_{0}^{*}$ as inputs for classification task. b. Comparison of test accuracy for the linear classifier trained on Bi-RNN outputs. The yellow curve represents accuracy with untrained (random) hidden neuron connections, the blue curve shows results from training with SCFF, the green curve shows Backprop results.
@@
After the first stage of unsupervised training, a linear classifier is trained on the hidden states from the final time step in both directions, as shown in the bottom of the Fig. 4a. The blue, orange and green curves in Fig. 4b depict the test accuracy of the linear output classifier with hidden connections trained using SCFF, with random (untrained) hidden connections, and with Backpropagation methods, respectively.
SCFF achieves a test accuracy of 90.33% $\pm$ 0.94% when trained with a one-layer Bi-RNN containing 500 hidden neurons in each direction (refer to Appendix F for further architectural details). It is below the performance of BackPropagation Though Time that easily reaches 96.00% $\pm$ 0.94% accuracy on this small task. However, SCFF avoids the issues of vanishing and exploding gradients of BPTT, as the gradients at each time step are calculated independently. This eliminates the dependency between time steps, providing a more stable training process which could be useful for future experiments on larger networks.
@@
Furthermore, the SCFF results are well above the model with untrained (random) input and hidden connections which plateaus at 80.78% $\pm$ 1.03% . Such model, in which only the output connections are trained, is akin to Reservoir Computing, a method that is often used to leverage physical systems on sequential data for neuromorphic applications [74]. SCFF provides a way to train these input and hidden layer connections in a simple, hardware-compatible way, and opens the path to a considerable gain of accuracy. This achievement opens the door for its extension to more complex tasks involving temporal sequences and its potential use in neuromorphic computing domains, such as dynamic vision sensors [75].
Overall, this result constitutes the first successful application of the FF approach to sequential data in an unsupervised manner. Future work could explore feedback mechanisms to enable top-down contextual modulation or extending SCFF with memory-augmented mechanisms, similar to Fast Weights [65], to improve temporal information retention.